### v2

# A Benchmark for Long-Form Medical Question Answering

Abstract
There is a lack of benchmarks for evaluating large language models (LLMs) in long-form medical question answering (QA). Most existing medical QA evaluation benchmarks focus on automatic metrics and multiple-choice questions. While valuable, these benchmarks fail to fully capture or assess the complexities of real-world clinical applications where LLMs are being deployed. Furthermore, existing studies on evaluating long-form answer generation in medical QA are primarily closed-source, lacking access to human medical expert annotations, which makes it difficult to reproduce results and enhance existing baselines. In this work, we introduce a new publicly available benchmark featuring real-world consumer medical questions with long-form answer evaluations annotated by medical doctors. We performed pairwise comparisons of responses from various open and closed-source medical and general-purpose LLMs based on criteria such as correctness, helpfulness, harmfulness, and bias. Additionally, we performed a comprehensive LLM-as-a-judge analysis to study the alignment between human judgments and LLMs. Our preliminary results highlight the strong potential of open LLMs in medical QA compared to leading closed models.
1 Introduction
The majority of existing LLM evaluation benchmarks in medical question answering (QA) have focused on automatic metrics and multiple-choice questions [14, 9, 19, 22]. Although valuable, such metrics and question formats fall short of reflecting the realistic settings of real-world clinical scenarios [24, 19] and do not fully assess or capture the nuances and factual accuracy of LLMs in the medical domain [9, 23, 25]. Additionally, there are concerns about the potential leakage of well-known benchmarks into the training data of LLMs that are subsequently evaluated on those same benchmarks [3]. Furthermore, benchmarks that have not leaked may contain label errors or be outdated [18], leading to flawed and unrealistic evaluations. Moreover, the limited research that has explored human evaluation of long-form medical QA has not made the associated labels publicly available, hindering reproducibility and the ability to study human annotations for insights to inform future work.
Figure 1: Overview of our benchmark creation process
To address these challenges, we introduce a new publicly available benchmark of real-world consumer medical questions with long-form answer evaluation, annotated by medical doctors. An overview of the process of building our benchmark is shown in Figure 1. Our contributions are as follows:
  • We develop a publicly available benchmark for long-form medical QA, based on real-world consumer health and medical questions.
  • We release the medical doctor annotations to the research community, along with our growing collection of real-world consumer medical questions.
  • We conduct a comprehensive analysis comparing human experts and LLM-as-a-judge for evaluating long-form answer generation in medical questions.
2 Data Preparation
To create a dataset of real-world consumer medical questions, we collected queries by users on our platform, Lavita Medical AI Assist. [^1] https://assist.lavita.ai/ In this phase, we collected all queries from 2023-10-31 (the date we launched our platform) to 2024-02-12 containing 4,271 inputs in 1,693 conversations. Conversations could be single-turn (1,011) or multi-turn (682). However, for now, we ignored the conversation types and processed queries independently, without considering whether they were part of the same conversation. The distribution of queries in conversations is shown in Figure 2. After collecting the queries, we deduplicated them, removed those from our sample question pool, [^2] By default, we show a few randomly selected questions on our website from a pool of trending medical questions. If a user’s question is from the pool, we remove it from our dataset. and filtered non-English entries using Lingua. [^3] https://github.com/pemistahl/lingua-py This step resulted in 2,698 queries.
Figure 2: The distribution of queries in conversations (excluding outliers)
2.1 Finding Medical Questions
User queries could be quite noisy and not all of them ask a clear and direct medical or health-related question. Additionally, some queries, even though medical, contain grammatical and/or spelling errors. Since manually filtering and correcting these cases is a time-consuming process, we prompted GPT-4444gpt-4-0125-preview to 1) tell if a user query contains a direct medical or health-related question, and 2) correct any grammatical or spelling errors in queries by retaining their original information and meaning. The full prompt template is shown in Figure 9.
2.1.1 Human Validation
Before relying on GPT-4’s annotations for medical question detection, we created a random sample of queries and had two human annotators label the same questions we asked GPT-4. We created a representative sample size of 337 based on Cochran‘s formula [2]. There is a 94% agreement among the two human annotators on whether a query contains a direct medical and health-related question. And, in those cases the two human annotators agreed, there’s also 91% agreement with GPT-4’s predictions. One common pattern we observed in disagreements between humans and GPT-4 was in cases in which it is hard to find a clear, direct, and independent question. For example, What dose of spermidine did they use? Even though this question is health-related and medical, it is not clear who they is referring to. These cases happen especially when a question is part of a multi-turn conversation which is not the focus of our study here. Considering the high agreement between human annotators and GPT-4, we separated all queries labeled by GPT-4 as a medical question.
2.1.2 Quality Check
To ensure the grammatically corrected version of queries by the model are not significantly different than the original ones, we computed [^5] We used the SequenceMatcher from the difflib library. the similarity between the original and corrected queries with an 85% threshold after manually inspecting some samples to determine a reasonable threshold. This process resulted in 1,446 queries identified as medical questions. It is worth noting that using GPT-4’s annotation, we were able to automatically cut 46% of queries that are non-medical. And, our prompt and pipeline could also be used in the future to filter out non-medical questions at a higher scale. Even though this is not the focus of our study, this experiment also shows the potential of using highly capable LLMs for medical question identification or classification.
To further check the quality of our questions, as the final step, one human annotator went through the 1,446 medical questions and removed any question that might have been wrongfully annotated by GPT-4 as a direct medical question. This process resulted in 1,298 remaining higher-quality medical questions.
2.2 Semantic Deduplication
In order to minimize similarity among the questions in our benchmark and to perform semantic deduplication, we conducted an intra-dataset similarity analysis to assess the degree of semantic similarity between the questions. For this analysis, we first computed the embeddings of all questions using OpenAI’s text-embedding-3-large model with a dimension size of 1024. We then constructed a distance matrix of the embeddings using cosine distance, defined as 1.0 minus the cosine similarity between two embedding vectors. Then we clustered embeddings based on DBSCAN [5] [^6] We used the scikit-learn implementation [17]. with:
[…]
We tried multiple thresholds and manually checked the clusters every time to ensure clusters were neither too specific nor too broad. This process resulted in 1,077 clusters. Then we selected the medoid of each cluster as the cluster’s representative. As a result, we ended up with 1,077 more semantically distinct questions.
2.3 Inter-Dataset Similarity Analysis
We also compared our dataset with three other well-known consumer health question datasets to measure the novelty of questions in our benchmark. These datasets include MedRedQA [15] with 51k pairs of consumer questions and their corresponding expert answers from Reddit (r/AskDocs/), HealthSearchQA [20] that includes 3,375 commonly searched consumer medical questions and introduced as a benchmark in the Med-PaLM project [20], and MASH-QA [28] with more than 34k question answer pairs sourced from WebMD. We separated a random sample of the same size as our benchmark from each of these datasets. To determine the similarity score between every two datasets, we used OpenAI’s text-embedding-3-large model to create 1024-dimensional embeddings for the questions. We then calculated the cosine similarity for all pairs of embeddings and took the average of these scores as the final similarity score between the datasets. The similarity score graph is shown in Figure 3. The similarity scores show a low overlap between our benchmark and existing datasets.
Figure 3: Similarity graph of consumer medical question answering datasets
2.4 Difficulty Level Annotation
Before creating our annotation data batches, to ensure that each batch is diverse and contains questions with various levels of sophistication, we annotated the difficulty level of questions by GPT-4. To define a scheme for difficulty levels of medical questions, we refined the medical tasks difficulty level scoring system introduced in Zhang et al., [26] and made it fit the medical question answering task. After consulting with medical doctors in the pre-annotation phase, we merged the original 5 categories into 3 since doctors found it challenging to distinguish the nuances among the original 5 categories. Our scheme is shown in Table 1. The prompt we designed for the annotation of difficulty levels based on our scheme can be found in Figure 11. Once difficulty levels are annotated, we are ready to set up our annotation task and create annotation batches.
2.5 Annotation Batches
We create annotation batches each with 100 questions. We randomly sample questions from each difficulty level category proportional to the distribution of questions in that category for each batch. This is to ensure that each batch has questions from all difficulty levels proportional to their distribution in the dataset. Next, using the selected models for evaluation (Section 3.1) for each batch, we generate answers for each question in each batch and randomize the order of the answers to prevent any association between the answer order and a specific model. The prompt for answer generation is shown in Figure 10, which is the same long-form question prompt in Singhal et al., 2023b [21] for consistency with prior work. Once batches and all answers are generated, we set our pairwise annotation task on LabelBox. [^7] https://labelbox.com/. Details of the annotation platform along with the annotation user interface can be found in Appendix A.
Table 1: Scheme of the difficulty levels of medical questions
Level Description
Basic Medical questions in this category are basic and straightforward. The answers become apparent immediately upon reading the question or can be easily located through a simple Google or Internet search. Some questions may require a minor application of real-world knowledge, rephrasing, or expanding on the information to find the answer.
Intermediate This category includes medical questions that are somewhat complicated, requiring a greater application of real-world knowledge. These questions tend to be detailed and may necessitate complex paraphrasing or simplification for clearer understanding. They can involve practical situations that require emotional support, psychological evaluations, and ethical considerations. Typical questions in this category might be similar to those found in USMLE exams. Furthermore, questions in this category might be based on vague symptom descriptions, making the diagnosis challenging, though they do not yet involve the most intricate scenarios of medical practice.
Advanced This category involves complex medical questions that require extensive and detailed medical knowledge. Questions at this level are often lengthy and intricate, relating to real-world scenarios that include actual medical cases with challenging diagnoses and treatments. The symptom descriptions can be highly vague, adding to the diagnostic challenge. These questions necessitate advanced multi-step thinking and decision-making, often involving new technologies, recent medical publications, or current global health issues like pandemics. This level demands a high level of decision-making skills, the ability to choose the best available option, and the demonstration of humane care, pushing the boundaries of medical expertise and ethical considerations.
3 Annotation and Evaluation
In building our benchmark, we aim to gain insights into the following research questions for long-form medical QA:
  • RQ1) How do smaller-scale open medical LLMs, trained on different vanilla models, perform compared to one another?
  • RQ2) What is the performance gap between open medical/non-medical models and closed state-of-the-art (SOTA) models?
  • RQ3) What is the effect of additional pretraining using medical data on relatively strong open vanilla models?
These insights could help us better understand the capabilities of open models and, ultimately, build improved open medical models that are preferred over commercial models due to the latter’s lack of transparency, privacy concerns, and unclear path toward HIPAA-compliant deployment.
3.1 Models
We selected the following set of open and closed medical and general-purpose LLMs for our evaluation: [^8] Details on model inference endpoints can be found in Appendix B AlpaCare [26], a series of LLMs instruction-tuned for medical tasks. Evaluations of these models have been conducted automatically on MedInstruct and iCliniq. The instruction tuning of AlpaCare was performed using MedInstruct-52k, a dataset built to cover diverse and high-quality examples for instruction tuning. BioMistral [10] is the first open-source biomedical model based on Mistral, further pre-trained on PubMed Central. BioMistral 7B has been tested on MMLU [6], MedQA [7], MedMCQA [16], and PubMedQA [8]. Additionally, the truthfulness of the model was evaluated on the medical subset of TruthfulQA [13]. BioMistral’s evaluations have been quantitative, with no human evaluation of the model’s long-form answer generation. For our evaluation, we selected the BioMistral 7B DARE model, as it achieved the highest average performance across all benchmarks reported in the BioMistral paper. We also selected two models from the Llama 3.1 family, including Llama-3.1-405B-Instruct and Llama-3.1-70B-Instruct [4]. We compare Llama-3.1-405B-Instruct, one of the most capable general-purpose open models, against GPT-4o, [^9] It’s worth noting that at the time we conducted the AlpaCare-13B vs. GPT-4 evaluation, GPT-4o had not been released. After the release of GPT-4o, we decided to replace GPT-4 with GPT-4o for our next evaluation batch to gain insights into this newer OpenAI model as well. one of the flagship commercial closed models. We also compare Llama-3.1-70B-Instruct against Meditron3-70B, a suite of LLMs specialized in clinical medicine, built on MEDITRON [1], with its base model being Llama-3.1-70B-Instruct.
3.2 Evaluation Criteria
There have been few studies that define a fine-grained annotation scheme for individual and pairwise evaluation of long-form answers in the medical QA domain. Perhaps the most comprehensive study is the series of works conducted in the development of the Med-PaLM models [20, 21].
Our initial draft of the annotation scheme was inspired by Med-PaLM’s approach. We merged some overlapping criteria in this scheme—for example, combining the "extent of possible harm" and "likelihood of possible harm" into a single criterion called harmfulness. We then independently shared our modified scheme with three medical doctors to gather their expert feedback. Our goal was to establish a distinct set of criteria that capture the most important aspects of long-form answer evaluation without being overly fine-grained. We specifically asked the doctors three questions: 1) Are all labeling criteria clear, or is there any confusion regarding any criterion? 2) Are there any additional overlapping criteria that could be merged? and 3) Are there any redundant or unnecessary criteria that could be removed? The doctors suggested further refinements, such as combining the axes of Unnecessary additional content and Missing important content into a single criterion called efficiency, which reflects how well an answer provides accurate medical knowledge without omitting relevant facts or including extraneous information. They also recommended rewording some criteria to make them more straightforward—for instance, changing "Which answer has a greater severity, extent, or likelihood of possible harm?" to "Which answer poses a higher risk of causing harm?". Our final annotation scheme is shown in Table 2.
3.3 Human Evaluation
We conducted our human evaluations with a group of three medical doctors, with two doctors assigned per batch, specializing in radiology and pathology. Before starting the main round of annotations, we shared our annotation scheme with the doctors and conducted a trial round with each on a small sample of questions. [^10] These trial questions were excluded from the main round of annotations. We then gathered the doctors’ feedback to ensure that all annotation criteria were clear and that there was no ambiguity regarding the instructions. After confirming clarity and receiving approval from the doctors, we proceeded with the main batches of annotations.
Table 2: Our pairwise comparison expert annotation scheme
Criterion Question Choices
Difficulty Classify the difficulty level of the given medical question as basic, intermediate, or advanced (please refer to the difficulty level guide for descriptions of each level). 1) Basic 2) Intermediate 3) Advanced
Correctness Which answer is more correct and accurate considering the current consensus of the scientific and clinical community? 1) Response A 2) Response B 3) Tie (equally good) 4) Neither (equally bad)
Helpfulness Which answer is more helpful? Measure helpfulness based on the model’s ability to assist users, considering the question’s intent.
Harmfulness Which answer poses a higher risk of causing harm?
Reasoning Which answer demonstrates better reasoning steps? (Correct rationale or manipulation of knowledge for answering the question.)
Efficiency Which answer provides accurate medical knowledge and descriptions without omitting important relevant facts or including extraneous information?
Bias Which answer contains information that is biased toward any demographic groups?
3.4 LLM-as-a-Judge
To design our LLM-as-a-judge prompt template we merged the pair-v2 system prompt in Zheng et al., [27], and the pairwise evaluation template in WildBench [12]. Our full prompt is shown in Figure 12. We make our prompt fit into the pairwise comparison of responses on a set of criteria instead of a single criterion. We ran two LLMs as our judges including gpt-4o-2024-08-06 and claude-3-5-sonnet-20241022.
3.4.1 Testing Robustness
Before using LLM judgments and comparing them with human evaluations, we conducted an analysis to assess the robustness and consistency of LLM judgments. Each LLM was run six times on each batch: three times with the default order of responses, Response A and B, and three times with the reversed order of responses. This was done to ensure that LLM judgments were not affected by positional biases of responses. We refer to the votes from the first three runs with the default order of responses as ab runs, and the votes from the three runs with the reversed order of responses as ba runs. Out of all 4,800 runs1111114 batches $\times$× 100 questions $\times$× 6 criteria per question $\times$× 2 LLM judges = 4,800 given by LLMs, there were 191 cases (%4) among ab runs and 213 cases (%4) in ba runs where the LLM was not consistent in its judgment. These inconsistencies in the gpt-4o model are approximately two times those in the claude model. Before moving to the next step, we resolved the inconsistencies by finding the majority vote for each run type.
After finding the majority votes, we compared the ab and ba run votes for each model individually. We found that in 811 cases (%17 of 4,800 total runs, 394 for gpt-4o and 417 for claude) there’s a disagreement in model judgments between ab and ba runs. In other words, when we changed the order of responses, in  %17 of cases, the models had a different judgment despite asking them to avoid positional biases. The highest disagreements were related to the efficiency and correctness criteria, and the least disagreement was for the bias criterion (details can be found in Appendix A.2.) Following Zheng et al., [27], we took a conservative approach and set the vote as tie and neither (for harmfulness and bias criteria) to resolve the disagreements. Then we compared gpt-4o and claude judgments and measured the correlation among their votes. There’s %71 percentage and %55 Cohen’s kappa agreement between the two models with chance agreement being %37. In the end, we resolve the disagreement between the two models based on the conservative approach described earlier and consider one final vote as our LLM-as-a-judge vote.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Medical DoctorsLLM-as-a-Judge
Figure 4: Distribution of votes by medical doctors (left) and LLM judge (right)
4 Results and Analysis
The distributions of the cumulative number of votes by medical doctors, as well as LLM-as-a-judge votes for the four batches, are shown in Figure 4. Starting with human evaluations on smaller-scale domain-specific models, we observe that AlpaCare-13B outperforms the smaller-scale model, BioMistral-7B. However, as expected, when compared to the much larger closed general-purpose model, GPT-4, AlpaCare-13B demonstrates lower performance across all criteria. The results become more interesting when comparing one of the flagship closed models, GPT-4o, with the state-of-the-art open model at the time, Llama-3.1 405B-Instruct, where Llama-3.1 outperforms GPT-4o across all aspects. This is especially promising for domains like healthcare, where user privacy is paramount, as deploying capable open models like Llama-3.1 could address privacy concerns by avoiding the need to send user data to third-party APIs with limited control. Finally, when comparing Meditron3-70B with its base vanilla model, Llama-3.1-70B-Instruct, we find that Meditron3-70B does not necessarily offer improvements over its base model. While we acknowledge that more rigorous and comprehensive evaluations are needed to generalize these conclusions, these results suggest that previous assumptions about the superiority of domain-specific medical or clinical models over general models or in-context learning approaches [11] may need to be revisited.
Turning to LLM-as-a-judge results and their comparison with human evaluations, we find general agreement across all batches and criteria. However, there remains a gap in the alignment between LLM votes and human labels.
4.1 Annotator Agreement
The annotator agreements for all batches across evaluation criteria are shown in Table 3. In each batch, and for each criterion, we calculated the percentage (observed) and chance agreements. To calculate the chance agreement, we first count the frequency of each vote option for both annotators in a batch. Then we calculate the marginal probabilities as the frequency of each voting category divided by the total number of samples in the batch. The chance agreement is the sum of the products of marginal probabilities for each category.
Table 3: Annotator agreement on annotation batches. P: Percentage (observed) agreement; C: Chance agreement
Batch 1 Batch 2 Batch 3 Batch 4
Criterion P C P C P C P C
Difficulty 0.57 0.42 0.5 0.39 0.4 0.34 0.37 0.33
Correctness 0.45 0.25 0.36 0.31 0.37 0.3 0.49 0.42
Helpfulness 0.47 0.27 0.42 0.33 0.23 0.19 0.2 0.19
Harmfulness 0.51 0.38 0.5 0.44 0.77 0.73 0.81 0.82
Reasoning 0.40 0.27 0.36 0.33 0.17 0.16 0.27 0.27
Efficiency 0.32 0.24 0.37 0.31 0.26 0.25 0.11 0.11
Bias 0.92 0.90 0.95 0.95 0.95 0.95 0.95 0.95
By looking at the results, we see that in general for the first two batches, there’s a slight to a fair level of agreement. However, for the last two batches, the agreement level is fairly low. To better understand the reason for the low agreement, we also took a look at individual judgments by annotators on each criterion for each batch. Figures 5, 6, 7, and 8 show the side-by-side comparison of votes in each batch separated by evaluation criteria.
Table 4 also shows the number of disagreement cases and the voting pairs associated with these disagreements. As seen in all batches, the most prominent source of disagreement occurs when annotators differ on whether one of the models is better or if a tie/neither verdict should be given. Simply put, these numbers suggest that in most disagreement cases, it may be more difficult for annotators to decide which response is better, and disagreements are less frequently related to completely opposing votes on which model is superior (Response A vs. Response B). This, in turn, highlights the challenging nature of making fine-grained judgments in long-form medical answer evaluations. These agreement numbers can serve as a baseline, and it would be interesting to explore in future work how greater alignment can be achieved among medical professionals when evaluating long-form medical question-answering.
Table 4: Annotator disagreement counts on annotation batches across evaluation criteria
Disagreement Count (%)
Vote pair Batch 1 Batch 2 Batch 3 Batch 4
Response A - Response B 29 (4.83%) 35 (5.83%) 42 (7%) 32 (5.33%)
Response A - Tie 115 (19.17%) 82 (13.67%) 196 (32.67%) 209 (34.83%)
Response B - Tie 47 (7.83%) 114 (19%) 24 (4%) 32 (5.33%)
Response A - Neither 43 (7.17%) 24 (4%) 14 (2.33%) 14 (2.33%)
Response B - Neither 27 (4.5%) 23 (3.83%) 17 (2.83%) 10 (1.67%)
Tie - Neither 32 (5.33%) 26 (4.33%) 32 (5.33%) 20 (3.33%)
5 Conclusion
In this work, we introduced a new publicly available benchmark with human expert annotations for long-form consumer medical question answering. Our preliminary results demonstrate the promising performance of open models compared to their closed commercial counterparts. Remarkably, open models, even without additional pretraining on medical domain data, perform on par with or even better than specialized models. We hope that by providing all medical expert labels, our benchmark can serve as a baseline for developing methods and guidelines to improve alignment among human experts in long-form medical QA, contributing to progress in this important task.
Appendix A Annotation Details
A.1 Label Analysis
Figures 5, 6, 7, and 8 show the comparison of votes by annotators in each batch separated by evaluation criteria.
Figure 5: Votes grouped by annotators - Batch 1
Figure 6: Votes grouped by annotators - Batch 2
Figure 7: Votes grouped by annotators - Batch 3
Figure 8: Votes grouped by annotators - Batch 4
A.2 LLM-as-a-Judge Statistics
Table 5 shows the number of disagreement cases across ab and ba runs for LLM judges. These numbers demonstrate how many times a model had a different judgment about an evaluation criterion for a question when we reversed the order of Response A and Response B. Both models had the highest inconsistency for the efficiency criterion.
Table 5: Count of disagreements between ab and ba runs by model and evaluation criteria
Criterion GPT-4o Claude 3.5 Sonnet
Correctness 72 76
Helpfulness 77 63
Harmfulness 54 38
Reasoning 88 52
Efficiency 103 181
Bias - 7
Appendix B Inference Endpoints
The details of the inference endpoints we used can be found in Table 6.
Table 6: Model endpoints for inference
Model name Inference endpoint Provider
AlpaCare-13B xz97/AlpaCare-llama2-13b Hugging Face
BioMistral 7B DARE BioMistral/BioMistral-7B-DARE Hugging Face
Llama-3.1-405B-Instruct meta-llama/Meta-Llama-3.1-405B-Instruct-Turbo Together AI
Meditron3-70B OpenMeditron/Meditron3-70B Hugging Face
Llama-3.1-70B-Instruct meta-llama/Meta-Llama-3.1-70B-Instruct Hugging Face
GPT-4 gpt-4-0125-preview OpenAI
GPT-4o gpt-4o-2024-05-13 OpenAI
Figure 9: The prompt template for annotating whether a query asks a direct medical or health-related question
Figure 10: The prompt template for generating answers for medical questions. There’s a minor difference between GPT-4/Claude’s prompt at the end to accommodate output response in JSON format.
Figure 11: The prompt template for annotating difficulty level of medical questions
Figure 12: LLM-as-a-judge prompt template